# Some properties of normal subgroups determined from character tables

### v1

@@ 1. Introduction
In general, if $N$ is a normal subgroup of a finite group $G$, the character table of $N$ can not be determined from the character table of $G$. In [6], the $G$-character table of $N$ obtained from the character table of $G$ is defined (see Section 2). They are non-singular matrices and provide certain arithmetical relations among significant integers associated to the irreducible characters of $N$ such as their character degrees, indices of inertia subgroups and ramification numbers. The character table of $G$ determines those irreducible characters of $G$ that are over the same minimal $G$-invariant character of $N$, which are defined as the sum of characters of $N$ belonging in the same orbit determined by the conjugation action of $G$ on the set of irreducible characters of $N$. These characters of $N$ form a basis of the $\mathbb{C}$-vector space of the $G$-invariant class functions of $N$, and they become relevant to $G$-character tables of $N$, playing a similar role that irreducible characters do to the character table of a group $G$.
@@ 3. Main results
We point out that it is not difficult to find examples that show that the reciprocal implication of the above result is not true (for instance, $G=D_{18}$ and $N=\Phi(G)$).
@@
It is not difficult to find examples satisfying the previous theorem. For instance, using GAP, the group $G=SmallGroup(120,3)$ has a cyclic normal subgroup $N$ of order 20 and we can check from Corollary 2.3$(B_{i})$, that $p=2$ divides $\widehat{\theta}(1)$ for every $\widehat{\theta}\in\operatorname{Min}_{G}(N)\setminus\operatorname{Lin}_{G}(N)$. Then, we deduce that $N$ has a normal $2$-complement by Theorem 3.10. In fact, the unique prime number dividing a degree of a minimal $G$-invariant character of $N$ is $2$. Therefore, a minimal $G$-invariant character of $N$ has either degree 1 (which means it is a linear $G$-invariant character of $N$) or $2$-power degree. However, in relation to Thompson’s theorem, $p=2$ does not divide every irreducible character degree of $N$ since $N$ is abelian.
@@
Let $G$ be a finite group and $\chi\in{\operatorname{Irr}}(G)$. We say that $(H_{\chi},\lambda_{\chi})$ is a linear character pair with respect to $\chi$ if $H_{\chi}$ is a normal subgroup of $G$, $\lambda_{\chi}\in\operatorname{Lin}_{G}(H_{\chi})$ and ${\operatorname{Irr}}(G|\lambda_{\chi})=\{\chi\}$ (so $\chi$ is m.i. character with notation in [12]).
Recall that a group $G$ is a monomial group or $M$-group if for every irreducible character $\chi\in{\operatorname{Irr}}(G)$ there exists a subgroup $H$ and a linear irreducible character $\varphi\in{\operatorname{Irr}}(H)$ such that $\chi=\varphi^{G}$. By Taketa’s theorem (theorem 5.12 and 5.13 in [11]), every $M$-group is solvable and the derived length of $G$ is bounded by the number of irreducible character degrees of $G$. Structural properties of $nM$-groups (and $sM$-groups) were introduced in [9] as a generalization of $M$-groups. A group $G$ is said to be a $nM$-group if for every irreducible character $\chi\in{\operatorname{Irr}}(G)$ there exists a normal subgroup $H$ and a linear irreducible character $\varphi\in{\operatorname{Irr}}(H)$ such that $\chi=\varphi^{G}$. Observe that in this situation, we have $\chi(1)=|G:H|$ and, using the Frobenius reciprocity, the ramification number of $\chi$ with respect to $N$ is equal to $1$.
@@
Trivially, every abelian group $G$ is an $nMI$-group considereing $\lambda_{\chi}^{G}=\chi$ for every irreducible character $\chi\in{\operatorname{Irr}}(G)$, and reciprocally. Moreover, the character table determines if a group $G$ is a $nMI$-group.
@@
A normal subgroup $N$ of a group $G$ is said to be a $G$-invariant $nMI$-subgroup if for every $\theta\in{\operatorname{Irr}}(N)$, there exists $(H_{\theta},\lambda_{\theta})$ a linear $G$-invariant character pai withr respect to $\theta$. In particular, when $N=G$ we have that $G$ is a $nMI$-group.
@@
$\operatorname{Mcd}_{G}(N)=\{\widehat{\theta}(1)\mid\widehat{\theta}\in\operatorname{Min}_{G}(N)\}=\{\widehat{\theta}_{i}(1)\mid i=1,\dots,k\}$
and by
$\operatorname{Lcd}_{G}(N)=\{\theta(1)\widehat{\theta}(1)\mid\theta\in{\operatorname{Irr}}(N)\}=\{\theta_{i}(1)\widehat{\theta}_{i}(1)\mid i=1,\dots,s\}.$
@@
$\Gamma^{i}_{G}(N)\subseteq\ker(\widehat{\theta_{i}})$
with $\theta_{i}(1)\widehat{\theta_{i}}(1)=f_{i}$, for $i\in\{1,\ldots,s\}$. In particular, $N$ is a hypercentral subgroup of $G$ and $\textbf{l}_{G}(N)\leq s=|\operatorname{Lcd}_{G}(N)|$.
Proof. We prove the theorem, using induction on $G$. Let $N_{0}$ be any normal subgroup of $G$ contained in $N$. We claim that $N/N_{0}$ satisfies the hypothesis of the theorem. Let $\theta\in{\operatorname{Irr}}(N/N_{0})$, then there exists a $G$-character pair with respect to $\theta$, say $(H_{\theta},\lambda_{\theta})$. Note that $\lambda_{\theta}\in{\rm Lin}_{G}(H_{\theta})$ and $\lambda_{\theta}^{N}=m\widehat{\theta}$, for some integer $m$. Hence,
$\ker\widehat{\theta}=\ker(\lambda_{\theta})^{N}=\bigcap\limits_{x\in N}(\ker\lambda_{\theta})^{x}=\bigcap\limits_{x\in N}\ker\lambda_{\theta}^{x}=\ker\lambda_{\theta}.$
Then $N_{0}\leq\ker\widehat{\theta}=\ker\lambda_{\theta}\leq H_{\theta}$, which means that $\lambda_{\theta}\in\operatorname{Lin}_{G}(H_{\theta}/N_{0})$ and so our claim is proved.
Suppose that $\ker\widehat{\theta}\not=1$ and we work on $\bar{G}=G/\ker\widehat{\theta}$ by adopting bar for the subgroups. If $\theta(1)\widehat{\theta}(1)=f_{i}$, for some $1\leq i\leq s$, then $\theta(1)\widehat{\theta}(1)$ is j-th smallest degree of ${\rm Lcd}_{\bar{G}}(\bar{N})$, for some $j\leq i$. By induction, we have
$\Gamma^{i}_{G}(N)\ker\widehat{\theta}/\ker\widehat{\theta}\leq\Gamma^{j}_{G}(N)\ker\widehat{\theta}/\ker\widehat{\theta}=\Gamma^{j}_{\bar{G}}(\bar{N})\leq\ker\widehat{\theta}/\ker\widehat{\theta}=1.$
Therefore, $\Gamma^{i}_{G}(N)\subseteq\ker(\widehat{\theta_{i}})$ as wanted.
We may assume $\widehat{\theta}$ is faithful. In this case, we claim that $H_{\theta}={\bf Z}(G)\cap N$. Since $\ker\lambda_{\theta}=\ker\lambda_{\theta}^{N}=\ker\widehat{\theta}=1$, then
$[H_{\theta},G]=\bigcap\limits_{\lambda\in{\rm Lin}_{G}(H_{\theta})}\ker\lambda\leq\ker\lambda_{\theta}=1.$
@@
This implies that $\theta(1)\widehat{\theta}(1)>\chi(1)\widehat{\chi}(1)$ for all $\chi\in{\operatorname{Irr}}(\tilde{N})$. Let $t=|{\rm Lcd}_{\tilde{G}}(\tilde{N})|$, it follows that $\theta(1)\widehat{\theta}(1)$ is k-th smallest number in $\operatorname{Lcd}_{G}(N)$, for some integer $k\geq t+1$. By induction, we have
$({\bf Z}(G)\cap N)\Gamma^{t}_{G}(N)/{\bf Z}(G)\cap N=\Gamma^{t}_{\tilde{G}}(\tilde{N})\leq\bigcap\limits_{\widehat{\chi}\in{\rm Min}_{\tilde{G}}(\tilde{N})}\ker\widehat{\chi}=({\bf Z}(G)\cap N)/({\bf Z}(G)\cap N).$
Hence, $\Gamma^{t+1}_{G}(N)=1=\ker\widehat{\theta}$ as wanted. So the proof is complete.
@@
If $G$ is a $nMI$-group and $1=\chi_{i}(1)<\chi_{2}(1)<\cdots<\chi_{n}(1)$ are the distinct degrees of irreducible characters of $G$, then
$\Gamma_{i}(G)\leq Ker(\chi_{i}),$
@@
Regarding the Corollary 3.19, we note that not all nilpotent groups verify that the class of nilpotency is bounded by the number of degrees of irreducible characters. For instance, using GAP, we can check that the group $G=SmallGroup(32,9)$ has nilpotency class 3 and only has two irreducible character degrees.
@@
Acknowledgment. This work was done during a visit of the first author at Universitat Politècnica de València(UPV) and Universitat de València(UV). She wishes to thank them for their hospitality. The results in this paper are part of the third author’s Ph.D. thesis, and he acknowledges the support of Instituto Tecnológico de Santo Domingo (INTEC) and Universidad Autónoma de Santo Domingo (UASD), Santo Domingo, (Dominican Republic).

### v2

@@ 1. Introduction
In general, if $N$ is a normal subgroup of a finite group $G$, the character table of $N$ can not be determined from the character table of $G$. In [6], the $G$-character table of $N$ obtained from the character table of $G$ is defined (see Section 2). They are non-singular matrices that provide certain arithmetical relations among significant integers associated to the irreducible characters of $N$ such as their character degrees, indices of inertia subgroups and ramification numbers. The character table of $G$ determines those irreducible characters of $G$ that are over the same minimal $G$-invariant character of $N$, which are defined as the sum of characters of $N$ belonging in the same orbit determined by the conjugation action of $G$ on the set of irreducible characters of $N$. These characters of $N$ form a basis of the $\mathbb{C}$-vector space of the $G$-invariant class functions of $N$, and they become relevant to $G$-character tables of $N$, playing a similar role that irreducible characters do to the character table of a group $G$.
@@ 3. Main results
We point out that it is not difficult to find examples that show that the reverse implication of the above result is not true (for instance, $G=D_{18}$ and $N=\Phi(G)$).
@@
It is not difficult to find examples satisfying the previous theorem. For instance, using GAP (see [7]), the group $G=SmallGroup(120,3)$ has a cyclic normal subgroup $N$ of order 20 and we can check from Corollary 2.3$(B_{i})$, that $p=2$ divides $\widehat{\theta}(1)$ for every $\widehat{\theta}\in\operatorname{Min}_{G}(N)\setminus\operatorname{Lin}_{G}(N)$. Then, we deduce that $N$ has a normal $2$-complement by Theorem 3.10. In fact, the unique prime number dividing a degree of a minimal $G$-invariant character of $N$ is $2$. Therefore, a minimal $G$-invariant character of $N$ has either degree 1 (which means it is a linear $G$-invariant character of $N$) or $2$-power degree. However, in relation to Thompson’s theorem, $p=2$ does not divide every irreducible character degree of $N$ since $N$ is abelian.
@@
Let $G$ be a finite group and $\chi\in{\operatorname{Irr}}(G)$. We say that $(H_{\chi},\lambda_{\chi})$ is a linear character pair with respect to $\chi$ if $H_{\chi}$ is a normal subgroup of $G$, $\lambda_{\chi}\in\operatorname{Lin}_{G}(H_{\chi})$ and ${\operatorname{Irr}}(G|\lambda_{\chi})=\{\chi\}$ (so $\chi$ is m.i. character with notation in [12]). By Problem (6.3) of [11], we have that $\chi$ and $\lambda_{\chi}$ are fully ramified with respect to $G/H_{\chi}$.
Recall that a group $G$ is a monomial group or $M$-group if for every irreducible character $\chi\in{\operatorname{Irr}}(G)$ there exists a subgroup $H$ and a linear irreducible character $\varphi\in{\operatorname{Irr}}(H)$ such that $\chi=\varphi^{G}$. By Taketa’s theorem (see Theorems 5.12 and 5.13 in [11]), every $M$-group is solvable and the derived length of $G$ is bounded by the number of irreducible character degrees of $G$. Structural properties of $nM$-groups were introduced in [9] as a generalization of $M$-groups. A group $G$ is said to be a $nM$-group if for every irreducible character $\chi\in{\operatorname{Irr}}(G)$ there exists a normal subgroup $H$ and a linear irreducible character $\varphi\in{\operatorname{Irr}}(H)$ such that $\chi=\varphi^{G}$. Observe that in this situation, we have $\chi(1)=|G:H|$ and, using the Frobenius reciprocity, the ramification number of $\chi$ with respect to $N$ is equal to $1$.
@@
Trivially, every abelian group $G$ is an $nMI$-group onsidering $\lambda_{\chi}^{G}=\chi$ for every irreducible character $\chi\in{\operatorname{Irr}}(G)$. Moreover, the character table determines if a group $G$ is a $nMI$-group.
@@
A normal subgroup $N$ of a group $G$ is said to be a $G$-invariant $nMI$-subgroup if for every $\theta\in{\operatorname{Irr}}(N)$, there exists $(H_{\theta},\lambda_{\theta})$ a linear $G$-invariant character pair with respect to $\theta$. In particular, when $N=G$ we have that $G$ is a $nMI$-group.
We remark that if $(H_{\theta},\lambda_{\theta})$ is a linear $G$-invariant character pair with respect to $\theta$, then not necessarily $(H_{\theta},\lambda_{\theta})$ is a linear character pair with respect to $\theta$ (see, for instance, Example 3.22)
@@
$\operatorname{Mcd}_{G}(N)=\{\widehat{\theta}(1)\mid\widehat{\theta}\in\operatorname{Min}_{G}(N)\}.$
Definition 3.18.
Let $N$ a normal subgroup of $G$. Then a leader $G$-character of $N$ is a character of $N$ defined by the product $\theta\widehat{\theta}$, for some $\theta\in{\operatorname{Irr}}(N)$. We denote by
$\operatorname{Lcd}_{G}(N)=\{\theta(1)\widehat{\theta}(1)\mid\theta\in{\operatorname{Irr}}(N)\}$
the set of degrees of leader $G$-characters of $N$.
Obverve that $|\operatorname{Mcd}_{G}(N)|$ and $|\operatorname{Lcd}_{G}(N)|$ may be different interger numbers (see Example 3.22). Moreover, by Corollary 2.3$(B_{i})$, degrees of the leader $G$-characters of $N$ are determined by the character table of $G$.
@@
$\Gamma^{i+1}_{G}(N)\subseteq\ker(\widehat{\theta_{i}})$
with $\theta_{i}(1)\widehat{\theta_{i}}(1)=f_{i}$ and $\theta_{i}\in{\operatorname{Irr}}(N)$, for $i\in\{1,\ldots,s\}$. In particular, $N$ is a hypercentral subgroup of $G$ and $\textbf{l}_{G}(N)\leq s=|\operatorname{Lcd}_{G}(N)|$.
Proof. We prove the theorem, using induction on $|G|$. Let $N_{0}$ be any normal subgroup of $G$ contained in $N$. We claim that $N/N_{0}$ satisfies the hypothesis of the theorem. Let $\theta\in{\operatorname{Irr}}(N/N_{0})$, then there exists a $G$-character pair with respect to $\theta$, say $(H_{\theta},\lambda_{\theta})$. Note that $\lambda_{\theta}\in{\rm Lin}_{G}(H_{\theta})$ and $\lambda_{\theta}^{N}=m\widehat{\theta}$, for some integer $m$. Hence,
$\ker(\widehat{\theta})=\ker(\lambda_{\theta}^{N})=\bigcap\limits_{x\in N}(\ker(\lambda_{\theta}))^{x}=\bigcap\limits_{x\in N}\ker(\lambda_{\theta}^{x})=\ker(\lambda_{\theta}).$
Then $N_{0}\leq\ker(\widehat{\theta})=\ker(\lambda_{\theta})\leq H_{\theta}$, which means that $\lambda_{\theta}\in\operatorname{Lin}_{G}(H_{\theta}/N_{0})$ and so our claim is proved.
Let $\theta\in{\operatorname{Irr}}(N)$. If $\theta(1)\widehat{\theta}(1)=1$, then $\theta\in\operatorname{Lin}_{G}(N)$ and trivially $\Gamma^{2}_{G}(N)\subseteq\ker(\theta)=\ker(\widehat{\theta})$. Suppose that $\theta(1)\widehat{\theta}(1)\not=1$. Suppose that $\ker(\widehat{\theta})\not=1$ and we work on $\bar{G}=G/\ker(\widehat{\theta})$ by adopting bar for the subgroups. If $\theta(1)\widehat{\theta}(1)=f_{i}$, for some $2\leq i\leq s$, then $\theta(1)\widehat{\theta}(1)$ is $j$-th smallest degree of ${\rm Lcd}_{\bar{G}}(\bar{N})$, for some $j\leq i$. By induction, we have
$\Gamma^{i+1}_{G}(N)\ker(\widehat{\theta})/\ker(\widehat{\theta})\leq\Gamma^{j+1}_{G}(N)\ker(\widehat{\theta})/\ker(\widehat{\theta})=\Gamma^{j+1}_{\bar{G}}(\bar{N})\leq\ker(\widehat{\theta})/\ker(\widehat{\theta})=1.$
Therefore, $\Gamma^{i+1}_{G}(N)\subseteq\ker(\widehat{\theta_{i}})$ as wanted.
We may assume $\widehat{\theta}$ is faithful. In this case, we claim that $H_{\theta}={\bf Z}(G)\cap N$. Since $\ker(\lambda_{\theta})=\ker(\lambda_{\theta}^{N})=\ker(\widehat{\theta})=1$, then
$[H_{\theta},G]=\bigcap\limits_{\lambda\in{\rm Lin}_{G}(H_{\theta})}\ker(\lambda)\leq\ker(\lambda_{\theta})=1.$
@@
This implies that $\theta(1)\widehat{\theta}(1)>\chi(1)\widehat{\chi}(1)$ for all $\chi\in{\operatorname{Irr}}(\tilde{N})$. Let $t=|{\rm Lcd}_{\tilde{G}}(\tilde{N})|$, it follows that $\theta(1)\widehat{\theta}(1)$ is $i$-th smallest number in $\operatorname{Lcd}_{G}(N)$, for some integer $i\geq t+1$. By induction, we have
$({\bf Z}(G)\cap N)\Gamma^{t+1}_{G}(N)/{\bf Z}(G)\cap N=\Gamma^{t+1}_{\tilde{G}}(\tilde{N})\leq\bigcap\limits_{\widehat{\chi}\in{\rm Min}_{\tilde{G}}(\tilde{N})}\ker(\widehat{\chi})=({\bf Z}(G)\cap N)/({\bf Z}(G)\cap N).$
Hence, $\Gamma^{t+2}_{G}(N)=1=\ker\widehat{\theta}$ as wanted. So the proof is complete.
@@
If $G$ is a $nMI$-group and $1=\chi_{1}(1)<\chi_{2}(1)<\cdots<\chi_{n}(1)$ are the distinct degrees of irreducible characters of $G$, then
$\Gamma_{i+1}(G)\leq\ker(\chi_{i}),$
@@
Regarding the Corollary 3.20, we note that not all nilpotent groups verify that the class of nilpotency is bounded by the number of degrees of irreducible characters. For instance, using GAP (see [7]), we can check that the group $G=SmallGroup(32,9)$ has nilpotency class 3 and only has two irreducible character degrees.
@@
Moreover, we can check that $N$ is not $nMI$-group since $N$ has an irreducible character $\theta$ such that $\theta(1)=2$ and it does not exist a linear pair $(H_{\theta},\lambda_{\theta})$ respect to $\theta$ such that $\theta$ and $\lambda_{\theta}$ are fully ramified.
Acknowledgment.
The authors would like to thank the referee for his or her comments which led to the improvement of this work. This work was done during a visit of the first author at Universitat Politècnica de València(UPV) and Universitat de València(UV). She wishes to thank them for their hospitality. The results in this paper are part of the third author’s Ph.D. thesis, and he acknowledges the support of Instituto Tecnológico de Santo Domingo (INTEC) and Universidad Autónoma de Santo Domingo (UASD), Santo Domingo, (Dominican Republic).